\begin{thebibliography}{10}

\bibitem{Aluffi} Aluffi, P., {\it Chern classes for singular hypersurfaces.} Trans. Amer. Math. Soc. (351) (1999), no. 10, 3989-4026.

\bibitem{Aluffi1} Aluffi, P., {\it Weighted Chern-Mather classes and Milnor classes of
hypersurfaces}, Adv. Stud. Pure Math. {\bf 29} (2000), 1--20.

\bibitem{Alu-Mar} Aluffi, P.; Marcolli, M., {\it
Feynman motives of banana graphs}.
[J] Commun. Number Theory Phys. 3, No. 1, 1-57 (2009).


\bibitem {Beh} Behrend, K.,
{\it Donaldson-Thomas type invariants via microlocal geometry}. Ann.
Math. (2) 170 (2009) No. 3, 1307-1338.

\bibitem{BS}  Brasselet,  J.-P. and  Schwartz, M.-H., {\it Sur les classes de Chern
d'un ensemble analytique complexe}.  Ast\'erisque {\bf 82-83} (1981), 93--147.

\bibitem{BLSS1}  Brasselet, J.-P.,  Lehmann, D., Seade, J.  and  Suwa, T.,
{\it Milnor numbers and classes of local complete intersections}. Proc. Japan
Acad. Ser. A Math. Sci. {\bf 75} (1999) No. 10, 179--183.

\bibitem{BLSS2}  Brasselet, J.-P.,  Lehmann, D., Seade, J.  and  Suwa, T.,
{\it Milnor classes of local complete intersections}. Transactions
A. M. S. {\bf 354} (2001), 1351--1371.

\bibitem{BSY}
Brasselet, J.-P. Sch\"urmann, J., Yokura, S., {\it
Hirzebruch classes and motivic Chern classes for singular spaces}.
J. Topol. Anal. {\bf 2} (2010), no. 1, 1-55.

\bibitem{BSS}   Brasselet, J.-P., Seade, J.  and  Suwa, T., {\it  Vector fields on singular varieties}.
Lecture Notes in Mathematics {\bf 1987}. Springer Verlag (2009).

\bibitem{BMS} Callejas-Bedregal, R., Morgado,  M. F. Z. and  Seade, J., {\it L\^e cycles and Milnor
classes} (Preprint - 2010)

\bibitem{CMSS} Cappell, S. E., Maxim, L. Sch\"urmann, J., Shaneson, J. L.
 {\it Characteristic classes of complex hypersurfaces}.
Adv. Math. {\bf 225} (2010), 2616-2647.

\bibitem{Dimca} Dimca,  A., {\it Sheaves in topology}. Universitext. Springer-Verlag,
Berlin, (2004).

\bibitem{Ful}  Fulton, W., {\it Intersection theory}. Ergebnisse der Mathematik und ihrer Grenzgebiete,  Springer-Verlag, Berlin, (1984).

\bibitem{GM}  Goresky, M. and  MacPherson, R., {\it Stratified Morse theory}. Ergeb. Math.
Grenzgeb. (3) 14, Berlin, Springer-Verlag, (1988).

 \bibitem{Massey0}  Massey, D. B., {\it The L\^e varieties. I.} Invent.
Math. {\bf 99}, n. 2, (1990),  357--376.

 \bibitem{Massey} Massey,  D. B., {\it L\^e Cycles and Hypersurface
Singularities}.  Lecture Notes in Mathematics {\bf 1615},
Springer-Verlag (1995).


\bibitem{MacP}  MacPherson, R., {\it Chern classes for singular algebraic varieties}. Annals of Math. {\bf 100}, n. 2, (1974), 423--432.

\bibitem{Max}  Maxim, L. {\it  On Milnor classes of complex hypersurfaces}. Topology of stratified spaces, 161–175, Math. Sci. Res. Inst. Publ., {\bf  58}, Cambridge Univ. Press, Cambridge, (2011).

\bibitem{MSS} Maxim, L.,   Saito, M.,  Sch\"urmann, J.
{\it Hirzebruch-Milnor classes of complete intersections}. Preprint 2012; arXiv:1205.6447v2 [math.AG].


\bibitem{O-Y} Ohmoto, T. and Yokura, S., {\it Product Formulas for
the Milnor Class}.  Bulletin of the Polish Acad. of Sciences Math.
Vol. 48, no. 4 (2000).

\bibitem{Par-Pr0}   Parusi\'nski, A. and Pragacz,  P., {\it A formula for the Euler characteristic of singular hypersurfaces}, J. Alg. Geom. {\bf 4} (1995), 337-351.

 \bibitem{Par-Pr}   Parusi\'nski, A. and Pragacz,  P., {\it
 Characteristic classes of hypersurfaces and characteristic
cycles}. J. Algebraic Geom. {\bf 10} (2001), 63--79.

\bibitem{PP}   Parusi\'nski, A. and Pragacz,  P., {\it
 Chern-Schwartz-Macpherson Classes and the Euler Characteristic of Degeneracy Loci and Special Divisors}.
 Journal of the American Math. Society, vol. 8, no. 4 (1995), 793--817.

\bibitem{Sabbah} Sabbah, C., {\it Quelques remarques sur la g\'eom\'etrie des espaces
conormaux.} Ast\'erisque {\bf 130} (1985), 161--192.

\bibitem{S} Sch\"{u}rmann, J.,
{\it Nearby cycles and characteristic classes of singular spaces}
ArXiv: 1003.2343v2. To appear in the proceedings of the fifth
Franco-Japanese Symposium on Singularities, Strasbourg 2009.

\bibitem{ST} Sch\"{u}rmann, J. and  Tib\u ar, M.,
{\it Index formula for MacPherson cycles of affine algebraic
varieties.}. Tohoku Math. J. {\bf 62} (2010), 29--44.

\bibitem{Sch}  Schwartz, M. H., {\it Classes caract\'eristiques d\'efinis par une stratification
 d'une vari\'et\'e analytique complexe}. C.R.\,Acad.\,Sci. Paris {\bf 260} (1965), 3262--3264 and 3535--3537.

\bibitem{ST} Sch\"{u}rmann, J. and  Tib\u ar, M.,
{\it Index formula for MacPherson cycles of affine algebraic
varieties.}. Tohoku Math. J. {\bf 62} (2010), 29--44.
\bibitem{Suwa} Suwa, T., {\it Classes de Chern des intersections
compl\`{e}tes locales}. C. R. Acad. Sci. Paris I Math., {\bf 324}
(1996), 67--70

\bibitem{SS}
Seade, J.  and   Suwa, T., {\it An adjunction formula for local complete
intersections}. Internat. J. Math. {\bf 9} (1998), 759--768.


\bibitem{Suwa}   Suwa, T.,
{\it Classes de Chern des intersections compl\`etes locales}.
C.R.\,Acad.\,Sci. Paris {\bf 324}  (1996),  67--70.


\bibitem{Yokura}  Yokura, S., {\it On characteristic classes of complete intersections}
Contemp. Math. {\bf 241} (1999),  349--369.

\end{thebibliography}

\documentclass[12pt]{article}
\usepackage{amsmath}
\usepackage{amssymb}
\usepackage{amsthm}
\usepackage{bbm}
\usepackage{paralist}
\usepackage[all]{xy}

\newtheorem{Thm}{Theorem}

\newtheorem{theorem}{Theorem}[section]
\newtheorem{lemma}[theorem]{Lemma}
\newtheorem{corollary}[theorem]{Corollary}
\newtheorem{proposition}[theorem]{Proposition}

\theoremstyle{definition}
\newtheorem{remark}[theorem]{Remark}
\newtheorem{definition}[theorem]{Definition}
\newenvironment{proaf}{{\noindent \sc Proof} }{\mbox{ }\hfill$\Box$
                        \vspace{1.5ex}
                        \par}

\def\C{\mathbb C}

\def\0{\underline 0}

\begin{document}

\title {\bf On the Milnor classes of local complete intersections\footnote{Research partially supported by CAPES, CNPq and FAPESP,
Brazil, and by CONACYT and PAPIIT-UNAM, Mexico. \newline $\quad$ {\it
Key-words:} Complete intersections, Milnor classes,
Whitney stratifications, Schwartz-MacPherson classes, Fulton-Johnson classes.
\newline   {\it Mathematics Subject Classification.} Primary ;
32S50, 32S15, 14B05  Secondary; 14C17, 14J17, 55S35  }}

\vspace{1cm}
\author{R. Callejas-Bedregal, M. F. Z. Morgado and J. Seade }

\date{\bf }
\maketitle

\begin{abstract}
In this work we study   algebraic, geometric and topological
properties of  the Milnor classes of local complete
intersections
 with arbitrary singularities. We describe first the
Milnor class of the intersection of a finite number of
hypersurfaces, under certain conditions of transversality, in
terms of the Milnor classes of the  hypersurfaces. Using this
description we obtain a Parusi\'{n}ski-Pragacz type formula, an
Aluffi type formula and a description of the Milnor class of the
local complete intersection in terms of the global L\^e cycles of
the hypersurfaces that define it. We consider next  the general
case of a local complete intersection $Z(s)$ defined by a regular
section $s$ of a rank $r$ holomorphic bundle $E$ over a compact
manifold $M$, $r \geq 2$. We notice that  $s$ determines a
hypersurface $Z(\tilde s)$ in the total space of the
projectivization $\mathbb{P}(E^{\vee})$ of the dual bundle
$E^{\vee}$, and we give a formula  expressing
 the total Milnor class of the local complete intersection $Z(s)$ in terms of the
 Milnor classes of the  hypersurface $Z(\tilde s)$.

\end{abstract}

\section*{Introduction}

If $Y$ is a singular variety  in a complex manifold $M$,   its total
Milnor class $\mathcal M_*(Y)$ is an element in its  Chow group
$A_*(Y)$ that measures the difference between its total
Schwarz-MacPherson class $c_*^{SM}(Y)$  and its total Fulton-Johnson
class   $c_*^{FJ}(Y)$  of $Y$. Both of these classes are
generalizations for singular varieties of the classical Chern
classes of manifolds. The total Milnor class actually has support in
the singular set ${\rm Sing}(Y)$, and there  is a Milnor class in
each dimension, from 0 to that of ${\rm Sing}(Y)$. In particular,
when $Y$ has only isolated singularities which are all local
complete intersections,  there is only a $0$-degree Milnor class,
which is an integer, and this is the sum of the local Milnor numbers
(by \cite{SS, Suwa}).

The concept of Milnor classes  appears first  implicitly in P.
Aluffi's work \cite{Aluffi, Aluffi1} on  $\mu$-classes for
hypersurfaces in algebraic manifolds. Milnor classes for
hypersurfaces also appear implicitly in A. Parusi\'nski and P.
Pragacz' article \cite {Par-Pr0}, though the actual name of Milnor
classes was coined later
  by various authors at about the same time (see \cite{BLSS1, BLSS2, Yokura, Par-Pr}).
  The case of local complete intersections was first envisaged in \cite {BLSS1, BLSS2}.

The study of Milnor classes is an active field of current research, with significant applications to other related areas (see for instance  \cite{Alu-Mar,Beh, BSY, Max}). Furthermore, this notion is  notably extended in \cite{CMSS} to that of
Milnor-Hirzebruch classes of locally complete intersections in algebraic manifolds. These  are  Hodge-theoretic classes that arise from the  Hirzebruch classes introduced by Brasselet {\it et al}  in \cite {BSY} (see also \cite{MSS}).

Milnor classes  are  mysterious ``objects"  that encode much
information about the varieties in question, and this is being
studied by various authors from several points of views. For
instance, when $Y$ is defined by a regular section of a very ample
holomorphic line bundle over $M$, it is  proved in  \cite{BMS} that
the Milnor classes determine the global L\^e cycles and {\it vice
versa}. The global L\^e cycles are  a natural extension of the local
L\^e cycles introduced by D. Massey in \cite{Massey0, Massey} for
holomorphic map-germs, which determine (among other things) the
topology of  the local Milnor fibre up to homeomorphism.

Yet, most of the work on Milnor classes in the literature  is for the case of hypersurfaces. Here we focus on the
 complete intersection case, which is much harder ({\it cf.}  \cite {BLSS1, BLSS2, MSS}).

We consider first a finite collection  $\{E_{i}\}$, $r\geq 2$,
 of holomorphic vector bundles $E_{i}$ over
$M$ of rank $d_{i}$. For each of these bundles we consider  a regular holomorphic section  $s_{i}:M\rightarrow E_{i}$
 and let $X_{i}$ be the $(n-d_{i})$-dimensional local
complete intersection  defined by the zeroes of
 $s_{i}$.  We assume  further that the   $X_i$ are equipped with  Whitney stratifications such that all the intersections amongst strata in the various $X_i$ are transversal.
Then  we prove a result (Theorem \ref{4.1}) that
describes the total Milnor class of $X$ in terms of the total Schwartz-MacPherson
and Milnor classes of  the $X_i.$
 Our starting point  is
the work by Ohmoto and Yokura in \cite{O-Y}, describing the total
Milnor class of  finite Cartesian products of hypersurfaces. As a consequence we get (Corollary \ref{11}):

\begin{Thm}
With the above hypothesis we have: $${\cal
M}(X)=(-1)^{r-1}c\left( \left( TM|_{X}\right)^{\oplus r-1}
\right)^{-1}\cap \displaystyle\sum_{i=1}^{r} a_{1,i}\cdot...\cdot
a_{r-1,i} \cdot {\cal M}(X_{i}),$$ where
$a_{j,i}=\left\{\begin{array}{ccl}
c^{Vir}(X_{j+1}) &\;{\rm if}\;& j\leq i\\
c^{SM}(X_{j}) &\;{\rm if}\;& j> i\\
\end{array}\right.$.
\end{Thm}

We focus next on the case when the bundles in question are all line bundles $L_i$, so each $X_i$ is a hypersurface. We get three types of applications:

\vspace{0,2cm}  \noindent
{\bf i)}
A Parusi\'nski-Pragacz type formula for local
complete intersections as above (Corollary \ref{P-P}).
This answers positively the expected description given
by Ohmoto and Yokura in \cite{O-Y} for the  total Milnor class of a local
complete intersection, as a polynomial in the Chern classes
$c_1(L_i),\;i=1,\dots,r.$ We  notice that a remarkable (different) generalization of the Parusi\'nski-Pragacz  formula for complete intersections has been given recently in \cite{MSS}.

\vspace{0,2cm}  \noindent
{\bf ii)}  A link between the Milnor classes of  a local complete intersection
$X$ as above and the global L\^e cycles of the hypersurfaces $X_i$ that define it (Theorem \ref{leformula}).

\vspace{0,2cm}  \noindent
{\bf iii)}  A   description of the total Milnor class of the local complete intersection $X$  in the vein of  Aluffi's formula in \cite{Aluffi} for hypersurfaces, using  Aluffi's $\mu$-classes (Corollary \ref{Aluffi}).

\vspace{0,2cm}

Finally we look at the
 general case: We consider  a holomorphic vector bundle $E$ of rank $r$ over
an $n$-dimensional compact complex
analytic manifold $M$,  and we let   $Z(s)$ be the zero set of a regular holomorphic
section $s$ of $E$. Then $Z(s)$  is an $(n-r)$-dimensional local complete
intersection. We look at the corresponding projectivized  bundle ${\mathbb P}(E^{\vee})$ and notice that the section $s$ induces
a section $\tilde s$ of the tautological bundle ${\mathcal O}_{{\mathbb P}(E^{\vee})}(1)$ over $\mathbb{P}(E^{\vee})$, and therefore it defines a hypersurface $Z(\tilde s)$ in $\mathbb{P}(E^{\vee})$. Then we give a formula expressing the total Milnor class of the local complete intersection $Z(s)$ in terms of the  total Milnor class of the hypersurface $Z(\tilde s)$  and the Chern classes of the various bundles in question. We prove (Theorem \ref{t:final}):

\begin{Thm} Let
$p:\mathbb{P}(E^{\vee})\rightarrow M$ be the projectivization of the
dual bundle $E^{\vee}$, so we have the tautological exact sequence
$$0\rightarrow F\rightarrow p^{*}E\rightarrow {\cal
O}_{\mathbb{P}(E^{\vee})}(1)\rightarrow 0.$$  Then, setting ${\cal
O}(1):={\cal O}_{{\mathbb P}(E^{\vee})} (1)$, we get:
$${\cal M}(Z(s)) = p_{*} \Big(\big[c\big(p^{*}(E^{\vee}) \otimes {\cal O} (1)\big)^{-1} \cdot c_{1}\big({\cal O}(1)\big)^{r-1} \cdot c(F)^{-1} \cdot c_{top}(F)\big] \cap {\cal M}(Z(\tilde{s}))\Big),$$
where $\tilde{s}$ is the section induced by $p^{*}s$ in ${\cal
O}(1)$.
\end{Thm}

We are grateful to Nivaldo Medeiros Jr., from
Universidade Federal Fluminense in Brazil, for fruitful
suggestions that helped us a lot to get our results in final form.
We are also grateful to Marcelo Jos\'e Saia, from USP at S\~ao
Carlos-Brazil, for helpful conversations.

\section{Derived Categories}

We  assume some basic knowledge on  derived categories,
hypercohomology and  sheaves of vanishing cycles as described in
\cite{Dimca}.

If $X$ is a complex analytic space then ${\cal D}^{b}_{c}(X)$
denotes the derived category of bounded, constructible complexes of
sheaves of $\C$-vector spaces on $X$. We denote the objects of
${\cal D}^{b}_{c}(X)$ by something of the form $F^{\bullet}$. The
shifted complex $F^{\bullet}[l]$ is defined by
$(F^{\bullet}[l])^{k}=F^{l+k}$ and  its differential is
$d^{k}_{[l]}=(-1)^{l}d^{k+l}$. The constant sheaf $\C_{X}$ on $X$
induces an object $\C_{X}^{\bullet} \in {\cal D}^{b}_{c}(X)$ by
letting $\C_{X}^{0}=\C_{X}$ and $\C_{X}^{k}=0$ for $k\neq 0$.

If $h:X\rightarrow \C$ is an analytic map and $F^{\bullet}\in
{\cal D}^{b}_{c}(X)$, then we denote the sheaf of vanishing cycles
of $F^{\bullet}$  with respect to $h$ by $\phi_{h}F^{\bullet}$.

For $F^{\bullet}\in {\cal D}^{b}_{c}(X)$ and $p \in X$, we denote
by ${\cal H}^{*}(F^{\bullet})_{p}$ the stalk cohomology of
$F^{\bullet}$ at $p$, and by $\chi(F^{\bullet})_{p}$ its Euler
characteristic. That is,
$$\chi(F^{\bullet})_{p}=\sum_{k}(-1)^{k}{\rm dim}_{\C}{\cal
H}^{k}(F^{\bullet})_{p}.$$
 We also denote by $\chi(X,F^{\bullet})$ the Euler characteristic
of $X$ with coefficients in $F^{\bullet}$, {\it i.e.},
$$\chi(X,F^{\bullet})=\sum_{k}(-1)^{k}{\rm dim}_{\C}\mathbb{H}^{k}(X,F^{\bullet}),$$
where $\mathbb{H}^{*}(X,F^{\bullet})$ denotes the hypercohomology
groups of $X$ with coefficients in $F^{\bullet}$.

When $F^{\bullet}\in {\cal D}^{b}_{c}(X)$ is ${\cal
S}$-constructible, where ${\cal S}$ is a Whitney stratification of
$X$, we  denote it by $F^{\bullet}\in {\cal D}^{b}_{{\cal S}}(X)$.
We have
\cite[Theorem 4.1.22]{Dimca}:
\begin{equation}\label{EulerCharact}\chi(X,F^{\bullet})=\sum_{S\in {\cal
S}}\chi(F^{\bullet}_{S})\chi(S),\end{equation} where
$\chi(F^{\bullet}_{S})=\chi(F^{\bullet})_{p}$ for an arbitrary point
$p \in S$.

For a subvariety $X$ in a complex manifold  $M$ we denote its conormal variety by
$T^{*}_{X}M$. That is, $$T^{*}_{X}M:={\rm closure}\;\{ (x, \theta) \in T^{*}M\;|\; x
\in X_{{\rm reg}}\;{\rm and}\; \theta_{|_{T_{x} X_{{\rm
reg}}}}\equiv 0\}\,,$$ where $T^{*}M$ is the cotangent bundle of $M$ and
$X_{{\rm reg}}$ is the regular part of $X$.

The following definition is standard in the literature:

\begin{definition}
Let $ X $ be an analytic subvariety of a complex manifold $ M $, $\{
S_{\alpha} \}$ a Whitney stratification of $M$ adapted to $ X $ and
$x\in S_\alpha$ a point in $X$. Consider $g:(M,x)\rightarrow (\C,0)$
a germ of holomorphic function such that $d_{x}g $ is a {\it
non-degenerate covector} at $x$ with respect to the fixed
stratification. That is, $d_{x}g \in T^{*}_{S_\alpha}M$ and $d_{x}g
\not\in T^{*}_{S^{'}}M$  for all stratum $S^{'} \neq S_\alpha$.  And
let $N$ be  a germ of a closed complex submanifold of $M$ which is
transversal to $S_\alpha$, with $N \cap S_\alpha=\{ x\}$. Define the
{\it complex link }  $l_{S_\alpha}$  of $S_\alpha$ by:
 \begin{center}$l_{S_\alpha}:= X\cap N \cap
B_{\delta}(x)\cap \{g=w\}\quad{\rm for}\;0<|w|<\!\!< \delta<\!\!<
1.$\end{center} The {\it normal Morse datum} of $S_\alpha$ is
defined by:
$$NMD(S_\alpha):=(X\cap N \cap B_{\delta}(x),l_{S_\alpha}),$$ and the {\it normal Morse index} $\eta(S_\alpha,F^{\bullet})$  of the stratum is:
\begin{center}$\eta(S_\alpha,F^{\bullet}):=\chi(NMD(S),F^{\bullet}),$\end{center}
where the right-hand-side means the Euler characteristic of the
relative hypercohomology. \end{definition}

By a result of M. Goresky and R. MacPherson in \cite[Theorem 2.3]{GM}  we
get that the number $\eta(S_\alpha,F^{\bullet})$ does not depends on
the choices of $x\in S_\alpha,\; g$ and $N$.

Notice that by \cite[Remark 2.4.5(ii)]{Dimca}, it follows that
\begin{equation}\label{RelativeEuler}\eta(S_\alpha,F^{\bullet})=\chi(X\cap N \cap
B_{\delta}(x),F^{\bullet})-\chi(l_{S_\alpha},F^{\bullet})\,.\end{equation}

\begin{remark}\label{ConstructibleFunction} Everything we have defined so far
for a constructible complex of sheaves is defined by J. Sch\"{u}rmann
and  M. Tib\u ar in \cite{ST} for constructible functions, and the two
 constructions are somehow equivalent. In fact, given
$F^{\bullet}\in {\cal D}^{b}_{c}(X)$, we have naturally associated
the  constructible function on $X$ given by
$$\beta(p)
=\chi(F^{\bullet})_{p}.$$ Moreover, by Sch\"{u}rmann \cite{S}, the
converse also holds, {\it i.e.}, given any \linebreak constructible
function $\beta$ on $X$ there is $F^{\bullet}\in {\cal
D}^{b}_{c}(X)$ such that
$$\beta(p) =\chi(F^{\bullet})_{p}.$$
\end{remark}

\section{Milnor classes}

\hspace{0.5cm} Let $M$ be an $n$-dimensional compact complex
analytic manifold and let $E$ be a holomorphic vector bundle over
$M$ of rank $d$. Let $Y$ be the zero set of a regular holomorphic
section of $E$, which is an $(n-d)$-dimensional local complete
intersection. Consider the virtual bundle $\tau(Y;M):= T
M|_{_{Y}}-E|_{_{Y}}$, where $T M$ denotes the tangent bundle of $
M$ and the difference is in the  KU-theory of $Y$. The virtual
homology class of $Y$ is defined by the Chern class of  $\tau(Y;M)$
via the Poincar\'e morphism, that is,
$$c^{Vir}(Y;M)=c(\tau(Y;M))\cap
[Y]:=(c^{}(T M|_Y)\cdot c^{}(E|_Y)^{-1})\cap [Y].$$
 Notice that restricted to
the regular part of $Y$, the bundle $E$ is isomorphic to the normal
bundle of $Y$ in $M$. Notice also that in the case we envisage here,
these classes coincide with the Fulton-Jhonson classes (see for
instance \cite{Ful}).

When there is no ambiguity, for simplicity we will denote the virtual bundle and the virtual  classes simply by $\tau(Y)$ and $c^{Vir}(Y)$.

On the other hand, consider the Nash blow up $\tilde{Y} \stackrel{\nu}{\rightarrow}
 Y$ of  $Y$, its Nash bundle ${\tilde T} \stackrel{\pi}{\rightarrow}
\tilde{ Y} $ and the Chern classes of $\tilde{T}$, $c^{j}(\tilde{T})
\in H^{2j}(\tilde{ Y})$, $j=1,\cdots,n$. The Mather classes of $Y$ are defined by
$$c^{Ma}_{k}( Y):=v_{*}(c^{n-d-k}(\tilde{T})\cap [\tilde{ Y}])\in H_{2k}( Y),\;\;k=0,\cdots,n \,.$$
We equip $Y$ with a Whitney stratification $Y_{\alpha}$.
The MacPherson classes are obtained from the Mather classes by considering
appropriate ``weights" for each stratum, determined by the local
Euler obstruction ${\rm Eu}_{{ Y_{\alpha}}}(x)$.  This is an integer associated   in \cite{MacP}
to each point $x \in Y_{\alpha}$.  It is proved \cite{MacP} that
there exists a unique set of integers $b_{\alpha}$ for which the
equation $\sum b_{\alpha} {\rm Eu}_{\bar{ Y}_{\alpha}}(x)=1$ is
satisfied for all points $x \in  Y$. Here, $\bar{ Y}_{\alpha}$ denotes the closure of the stratum, which is itself analytic and therefore it has its own local
Euler obstruction ${\rm Eu}_{\bar{ Y}_{\alpha}}$ and Mather classes, and the sum runs over all
strata $ Y_{\alpha}$ containing $x$ in their closure.

Then the  MacPherson class of degree $k
$ is defined by
$$c^{M}_{k}( Y):=\sum
b_{\alpha}\;i_{*}(c^{Ma}_{k}(\bar{ Y}_{\alpha})),$$
where  and
$i:\bar{ Y}_{\alpha}\hookrightarrow  Y$ is the inclusion map.

We remark that by \cite{BS}, the MacPherson classes coincide,
up to Alexander duality, with the classes defined previously by
M.-H. Schwartz in \cite{Sch}. Thus, following the modern
literature (see for instance \cite{Par-Pr, BLSS2, BSS}), we call
these the  Schwartz-MacPherson classes of $Y$ and  denote them
 by $c^{SM}_{k}(Y)$.

\begin{definition}
The Milnor class of $Y$ is:
$${\cal M}(Y):=(-1)^{n-d}\left(c^{Vir}(Y)-c^{SM}(Y) \right).$$
\end{definition}

\section{Milnor classes and the diagonal embedding}

\vspace{0,3cm}  Let $M$ be  as before, an $n$-dimensional
compact complex analytic manifold, and set $M^{(r)}:=M\times \cdots \times M$. In this section we let $E$ be a holomorphic vector
bundle over  of rank $d$. Consider
$\Delta: M\rightarrow M^{(r)}$ the diagonal morphism, which is a
regular embedding of codimension $nr-n$. Following \cite[Chapter
6]{Ful} we have that $\Delta$ induces the Gysin homomorphism
$$\Delta^{!}:A_{k}(M^{(r)})\rightarrow A_{k-nr+n}(M),$$ given by
$\Delta^{!}({\alpha}_1\times\cdots \times
{\alpha}_r)={\alpha}_1\cdots {\alpha}_r,$ where the right-hand side
is the intersection product of cycles as defined in \cite[Chapter
8]{Ful}. Topologically this homomorphism can be described by
capping with the orientation class in $H^{2(r-1)n}(M^{(r)},M^{(r)}-M)$ given
by the regular embedding $\Delta: M\rightarrow M^{(r)}.$

Let $t$ be a regular holomorphic section of $E$. Hence the set of
the zeros of $t$, $Z(t)$, is a closed subvariety of $M^{(r)}$ of
dimension $nr-d$.

\begin{proposition}\label{t:1} The Gysin morphism satisfies: $$\Delta^{!}\left( \;c^{Vir}(Z(t))\;\right)=
c\left( \left( TM|_{Z(\Delta^{*}t)}\right)^{\oplus r-1} \right)\cap c^{Vir}(Z(\Delta^{*}t))\,.$$\end{proposition}

\begin{proof}
By definition of the virtual class we have $$\Delta^{!}\;c^{Vir}(Z(t))=\Delta^{!}\left(
c\left( TM^{(r)}|_{Z(t)}\right)\cdot c\left(
{E}|_{Z(t)}\right)^{-1} \cap [Z(t)] \right).
$$
Hence, by \cite[Proposition 6.3]{Ful}, we have that
$$\Delta^{!}\;c^{Vir}(Z(t))=c\left(\Delta^{*}\left(
TM^{(r)}|_{Z(t)}\right)\right)\cdot c\left(
\Delta^{*}\left({E}|_{Z(t)}\right)\right)^{-1} \cap
\Delta^{!}\;[Z(t)].$$
Note that
$\Delta^{*}\left({E}|_{Z(t)}\right)=\Delta^{*}{E}|_{Z(\Delta^{*}t)}$
and, by \cite[Proposition 14.1]{Ful}, $\Delta^{!}\;[Z(t)]=
[Z(\Delta^{*}t)]$.
Moreover, since $\Delta^{*} TM^{(r)}= TM\oplus \cdots \oplus TM$, we
have $$c\left(\Delta^{*}\left(
TM^{(r)}|_{Z(t)}\right)\right)=c\left(\left(
TM|_{Z(\Delta^{*}t)}\right)^{\oplus\;r}\right).$$

Then,
\begin{center}$\left.\begin{array}{rcl}\Delta^{!}\;c^{Vir}(Z(t))&=&c\left(\left(
TM|_{Z(\Delta^{*}t)}\right)^{\oplus\;r}\right)\cdot
c\left(\Delta^{*}E|_{Z(\Delta^{*}t)}\right)^{-1}\cap \;[Z(\Delta^{*}t)]\\
~&=& c\left( \left( TM|_{Z(\Delta^{*}t)}\right)^{\oplus r-1}
\right)\cap c^{Vir}(Z(\Delta^{*}t)).\\
\end{array}\right.$
\end{center}
\end{proof}

For a subvariety $X$ of $M$, we denote its conormal variety by
$T^{*}_{X}M$, that is, $T^{*}_{X}M:={\rm closure}\;\{ (x, \theta) \in T^{*}M\;|\; x
\in X_{{\rm reg}}\;{\rm and}\; \theta|_{_{T_{x} X_{{\rm
reg}}}}\equiv 0\}$, where $T^{*}M$ is the cotangent space of $M$ and
$X_{{\rm reg}}$ is the regular part of $X$. Let $L( M)$ be the free abelian group of all cycles generated by the
conormal spaces $T^{*}_{X} M$, where $X$ varies over all subvarieties of $ M$.

Let  $\{ S_{\alpha} \}$ be a Whitney stratification of $M$
adapted to $ X $ and $x\in S_\alpha$ a point in $X$. Let $F(M)$ be the free
abelian group of constructible functions on
$M$.
Define the function
$Ch: F( M)\rightarrow L( M)$ by:
$$Ch(\xi):=\sum_{\alpha} (-1)^{\dim S_\alpha}\eta(S_\alpha,\xi)\cdot  T^{*}_{\overline{S}_\alpha} M.$$

Now consider the projectivized cotangent bundles $\mathbb{P}(T^{*}M)$ and  $\mathbb{P}(T^{*}(M^{(r)})$; we set $\mathbb{P}(T^{*}M\oplus \cdots \oplus T^{*}M):=\mathbb{P}((T^{*}M)^{\oplus r})$. Notice that one has a fibre square
 diagram (see \cite[pag. 428]{Ful}):
$$
\xymatrix {    \mathbb{P}((T^{*}M)^{\oplus r})
\ar[r]^{\delta} \ar[d]_{p} & \mathbb{P}(T^{*}(M^{(r)}))  \ar[d]^{\pi^{(r)}} \\ M \ar[r]^{\Delta}  & M^{(r)}  }
$$
where $\pi^{(r)}$ is the natural proper map. Let $i: \mathbb{P}(T^{*}M) \to \mathbb{P}((T^{*}M)^{\oplus r}) $ be the morphism
 induced by the diagonal embedding
 $T^{*}M \to T^{*}M \oplus \cdots \oplus T^{*}M$.

\begin{lemma}\label{L:4} Let $Z(t)$ be as in Proposition \ref{t:1}. Assume that $Z(t)$ admits a Whitney stratification $\{S_\alpha\}$ transversal to
$\Delta(M)$. Then:
$$\delta^{!}\;[\mathbb{P}(Ch({\mathbbm 1}_{Z(t)}))]=(-1)^{nr-n}\; i_{*}\;[\mathbb{P}(Ch({\mathbbm 1}_{Z(\Delta^{*}t)}))] \,,$$
where ${\mathbbm 1}_{(\,)}$ denotes the characteristic function.
\end{lemma}

\begin{proof} Since the stratification $\{S_{\alpha}\}$  is transversal to $\Delta(M)$, we have that
$\{\Delta^{-1}(S_{\alpha})\}$ is a Whitney stratification of
$Z(\Delta^{*}t)$.  By definition, $$\mathbb{P}(Ch({\mathbbm 1}_{Z(t)}))=\sum
m_{\alpha}\mathbb{P}\left(\overline{T_{S_{\alpha}}^{*}M^{(r)}}\right),$$
 where
$m_{\alpha}:= (-1)^{nr-d-1}\chi\left(\phi_{f|Z(t)}F^{\bullet}
\right)_{z}$, with $z \in S_{\alpha}$, $F^{\bullet}$ the complex sheaf
defined by $\chi( F^{\bullet})_{p}={\mathbbm 1}_{Z(t)}(p)$, a germ $f:
(M^{(r)},z)\rightarrow (\C,0)$ such that satisfies $(z,d_{z}f) \not\in
\overline{T_{S_{j}}^{*}M^{(r)}}$ for all strata $S_{\beta} \neq S_{\alpha}$ and
$\phi_{f|Z(t)}F^{\bullet}$ is the sheaf of  vanishing cycles of $f$ restricted to $Z(t)$.
Analogously,
$$\mathbb{P}(Ch({\mathbbm 1}_{Z(\Delta^{*}t)}))=\sum
n_{\alpha}\mathbb{P}\left(\overline{T_{\Delta^{-1}(S_\alpha)}^{*}M}\right),$$
where $n_{\alpha}:=
(-1)^{n-d-1}\chi\left(\phi_{g|Z(\Delta^{*}t)}G^{\bullet}
\right)_{x}$, with $x \in \Delta^{-1}(S_{\alpha})$, $G^{\bullet}$ the
complex sheaf defined by $\chi(
G^{\bullet})_{q}={\mathbbm 1}_{Z(\Delta^{*}t)}(q)$ and a germ $g:
M,x\rightarrow \C,0$ such that satisfies $(x,d_{x}g) \not\in
\overline{T_{\Delta^{-1}(S_{\beta})}^{*}M}$ for all strata $S_{\beta} \neq
S_{\alpha}$.

\vspace{0.2cm}These definitions do not depend on the choices of $
(z,f) $ and $(x,g) $ respectively. So we fix $ x $ and $ g $ in the
second definition, and take $ z = (x,\cdots,x) $ and $ f $ such that
$\Delta^{*}f=g$.

Note that, since $\Delta^{*}F^{\bullet}=G^{\bullet}$ and
$$\phi_{\Delta^{*}(f|Z(t))}\Delta^{*}F^{\bullet}=\Delta^{*}\left(\phi_{f|Z(t)}F^{\bullet}\right)\,,$$
we have: $$\chi\left(\phi_{g|Z(\Delta^{*}t)}G^{\bullet}
\right)_{x}=\chi\left(\phi_{f|Z(t)}F^{\bullet}
\right)_{x,\cdots,x}\,.$$ Hence
\begin{equation}\label{6}m_{\alpha}=(-1)^{nr-n} n_{\alpha}. \end{equation}
Notice that
\begin{equation}\label{7}\delta^{!}\;[\mathbb{P}\left(\overline{T_{S_{i}}^{*}M^{(r)}}\right)]=
i_{*}\;[\mathbb{P}\left(\overline{T_{\Delta^{-1}(S_{i})}^{*}M}\right)].\end{equation}
Therefore, by equations (\ref{6}) and (\ref{7}),
$$\delta^{!}\;[\mathbb{P}(Ch({\mathbbm 1}_{Z(t)}))]=(-1)^{nr-n}\;
i_{*}\;[\mathbb{P}(Ch({\mathbbm 1}_{Z(\Delta^{*}t)}))].$$\end{proof}

\begin{theorem}\label{T:2} With the assumptions of Lemma \ref{L:4} we have: $$\Delta^{!}\left( \;c^{SM}(Z(t))\;\right)=
c\left( \left( TM|_{Z(\Delta^{*}t)}\right)^{\oplus r-1} \right)\cap c^{SM}(Z(\Delta^{*}t))\,.$$\end{theorem}

\begin{proof} By the description of Schwartz-MacPherson classes due to C.
Sabbah in \cite{Sabbah}, we have that
$$c^{SM}(Z(t))=(-1)^{nr-1}c\left( TM^{(r)}|_{Z(t)} \right) \cap
\pi_{*}^{(r)} \left( c({\cal O}_{r}(1)^{-1})\cap
[\mathbb{P}(Ch({\mathbbm 1}_{Z(t)}))]\right),$$
where ${\cal O}_{r}(1)$ denotes
the tautological line bundle on the projectivisation
$\mathbb{P}(T^{*}M^{(r)})\rightarrow M^{(r)}$.
Using  again \cite[Proposition 6.3]{Ful} we get:
\begin{equation}\label{8}{\small \Delta^{!}\;c^{SM}(Z(t))=(-1)^{nr-1}c\left(\Delta^{*}\left(
TM^{(r)}|_{Z(t)} \right)\right) \cap \Delta^{!}\pi_{*}^{(r)} \left(
c({\cal O}_{r}(1))^{-1})\cap
[\mathbb{P}(Ch({\mathbbm 1}_{Z(t)}))]\right).}\end{equation}
And by \cite[Theorem 6.2]{Ful} we have,
\begin{equation}\label{9}\Delta^{!}\pi_{*}^{(r)} \left( c({\cal
O}_{r}(1))^{-1})\cap [\mathbb{P}(Ch({\mathbbm 1}_{Z(t)}))]\right)= p_{*} \left(
c(\delta^{*}{\cal O}_{r}(1))^{-1})\cap
\delta^{!}\;[\mathbb{P}(Ch({\mathbbm 1}_{Z(t)}))]\right).\end{equation}
Since $\delta^{*}{\cal O}_{r}(1)= {\cal
O}_{\mathbb{P}\left((T^{*}M)^{\oplus r} \right)}(1)$ is the
tautological line bundle on the projectivisation
$\mathbb{P}((T^{*}M^{(r)})^{\oplus r})\rightarrow M^{(r)}$, by Lemma \ref{L:4} and
the equations (\ref{8}) and (\ref{9}), we get:
$$\begin{array}{l}
\Delta^{!}\left( \;c^{SM}(Z(t))\;\right)= (-1)^{n-1}c\left( \left(
TM|_{Z(\Delta^{*}t)}\right)^{\oplus r} \right)\cap\\\quad \quad\quad\quad\quad\quad\quad \,\;\cap\; p_{*} \left(
c({\cal O}_{\mathbb{P}\left((T^{*}M)^{\oplus r}
\right)}(1))^{-1}\cap
i_{*}\;[\mathbb{P}(Ch({\mathbbm 1}_{Z(\Delta^{*}t)}))]\right).\end{array}$$
Using again the description of Schwartz-MacPherson due to Sabbah ({\it loc. cit.}) we obtain:
$$\Delta^{!}\left( \;c^{SM}(Z(t))\;\right)=c\left( \left( TM|_{Z(\Delta^{*}t)}\right)^{\oplus r-1}
\right)\cap c^{SM}(Z(\Delta^{*}t)).$$

\end{proof}

\begin{corollary} \label{10}With the previous assumptions we have: $$\Delta^{!}{\cal M}(Z(t))=(-1)^{nr-n}c\left( \left( TM|_{Z(\Delta^{*}t)}\right)^{\oplus
r-1} \right)\cap{\cal M}(Z(\Delta^{*}t))\,.$$\end{corollary}

\begin{proof}
Using Proposition  \ref{t:1} and Theorem \ref{T:2} , we have:
\begin{center}$\left.\begin{array}{rcl}
\Delta^{!}{\cal M}(Z(t))&=&\Delta^{!}\left(
(-1)^{nr-d}(c^{Vir}(Z(t))- c^{SM}(Z(t))) \right)\\
~&=&(-1)^{nr-d}c\left( \left( TM|_{Z(\Delta^{*}t)}\right)^{\oplus
r-1} \right)(c^{Vir}(Z(\Delta^{*}t))- c^{SM}(Z(\Delta^{*}t)))\\
~&=&(-1)^{nr-n}c\left( \left( TM|_{Z(\Delta^{*}t)}\right)^{\oplus
r-1} \right)\cap {\cal M}(Z(\Delta^{*}t)).\\
\end{array}\right.$
\end{center}\end{proof}

\section{Intersection product formula}

\hspace{0.5cm} As before, let $M$ be an $n$-dimensional compact complex analytic
manifold. Let $\{E_{i}\}$, be a finite collection of holomorphic vector bundles  over
$M$ of rank $d_{i}$, $2\leq i\leq r$. For each of these bundles, let  $s_{i}:M\rightarrow E_{i}$ be a regular holomorphic section
 and $X_{i}$ the $(n-d_{i})$-dimensional local
complete intersections  defined by the zeroes of
 $s_{i}$. In this section we assume that the   $X_i$ are equipped with  Whitney stratifications such that all the intersections amongst strata in the various $X_i$ are transversal.

 Let $p_i:M^{(r)}\rightarrow M$ be the $i^{th}$-projection,
then we have the holomorphic exterior product section
$$s=s_1\oplus\cdots\oplus s_r:M^{(r)}\rightarrow
p_1^*E_1\oplus\cdots \oplus p_r^*E_r,$$ given by
$s(x_1,\dots,x_r)=(s_1(x_1),\dots, s_r(x_r)).$ Then
$Z(s)=X_1\times\cdots \times X_r$ and $Z(\Delta^*(s))=X_{1}\cap
\cdots \cap X_{r}.$ Set $X=Z(\Delta^*(s)).$

The next result describes the total Milnor class of $X$ in terms
of the total Schwartz-MacPherson and Milnor classes of  the $X_i.$

\begin{theorem} \label{4.1} In the above conditions we have:
 $${\cal M}(X)=(-1)^{nr-n}c\left( \left(
TM|_{X}\right)^{\oplus r-1} \right)^{-1}\cap
\sum\;(-1)^{(n-d_{1})\epsilon_{1}+\cdots+(n-d_{r})\epsilon_{r}}
 P_{1}\;\cdot...\cdot\; P_{r}\in A_{*}(X),$$
where the sum runs over all choices  of $P_{i}\in \left\{{\cal
M}(X_i),c^{SM}(X_i)\right\},\,i=1,\dots,r,$  except
$(P_{1},\cdots,P_{r})=(c^{SM}(X_1),\cdots,c^{SM}(X_r))$ and where
$$\epsilon_{i}=\left\{\begin{array}{rcl} 1&,& \mbox{if} \;P_{i}=c^{SM}(X_i)\\
0&,& if \;P_{i}={\cal M}(X_i)\\
\end{array}\right. .$$

\end{theorem}

Thus for instance, when
$r=2$ we have:
$${\cal M}(X)=c\left( \left( TM|_{X}\right) \right)^{-1}\cap
\qquad \qquad\qquad\qquad   \qquad \qquad\qquad\qquad  \qquad
\qquad\qquad\qquad  $$
$$\quad \Big((-1)^{n}{\cal
M}(X_1)\cdot {\cal M}(X_2)+ (-1)^{d_{1}} c^{SM}(X_1)\cdot  {\cal
M}(X_2)+(-1)^{d_{2}}{\cal M}(X_1)\cdot  c^{SM}(X_2)\Big).$$ For
$r= 3$ we get: {\small
$${\cal M}(X)= c\left( \left( TM|_{X}\right)^{\oplus 2}
\right)^{-1}\cap \;\; \Big({\cal M}(X_1)\cdot {\cal M}(X_2)\cdot
{\cal M}(X_3)+ (-1)^{(d_{1}+d_{2})} c^{SM}(X_1)\cdot
c^{SM}(X_2)\cdot  {\cal M}(X_3) +$$
$$+(-1)^{(d_{1}+d_{3})} c^{SM}(X_1)\cdot {\cal
M}(X_2)\cdot  c^{SM}(X_3)+(-1)^{(d_{2}+d_{3})} {\cal M}(X_1)\cdot
c^{SM}(X_2)\cdot  c^{SM}(X_3)+$$
$$+(-1)^{(n-d_{1})} c^{SM}(X_1)\cdot {\cal M}(X_2)\cdot {\cal
M}(X_3)+(-1)^{(n-d_{2})} {\cal M}(X_1)\cdot c^{SM}(X_2)\cdot {\cal
M}(X_3)+$$
$$+(-1)^{(n-d_{3})} {\cal M}(X_1)\cdot {\cal M}(X_2)\cdot  c^{SM}(X_3)
\Big),$$} and so on.

\vskip.2cm

\begin{proaf}{of Theorem \ref{4.1}.}
By Corollary \ref{10},
$$\Delta^{!}{\cal M}(Z(s))=(-1)^{nr-n}c\left( \left( TM|_{Z(\Delta^{*}s)}\right)^{\oplus
r-1} \right)\cap{\cal M}(Z(\Delta^{*}s)).$$ Thus,
$${\cal M}(X)=(-1)^{nr-n}c\left( \left(
TM|_{Z(\Delta^{*}s)}\right)^{\oplus r-1}
\right)^{-1}\cap\Delta^{!}{\cal M}(X_{1}\times \cdots \times
X_{r}),$$
 and using the description of the Milnor classes of a product due to   \cite[Theorem 3.3]{O-Y}, we have:
$${\cal M}(X)=(-1)^{nr-n}c\left( \left(
TM|_{X}\right)^{\oplus r-1} \right)^{-1}\cap
\;\sum(-1)^{(n-d_{1})\epsilon_{1}+\cdots+(n-d_{r})\epsilon_{r}}\Delta^{!}\left(
 P_{1}\times ...\times  P_{r}\right),$$
where the sum runs over all choices  of $P_{i}\in \left\{{\cal
M}(X_i),c^{SM}(X_i)\right\},\,i=1,\dots,r,$  except
$(P_{1},\cdots,P_{r})=(c^{SM}(X_1),\cdots,c^{SM}(X_r))$ and
where $$\epsilon_{i}=\left\{\begin{array}{rcl} 1&,& if \; P_{i}=c^{SM}(X_i)\\
0&,& if \; P_{i}={\cal M}(X_i)\\
\end{array}\right. .$$
Therefore the result follows because  $$\Delta^{!}\left(
P_{1}\times ...\times P_{r}\right)= P_{1}\;\cdot...\cdot\;
P_{r}\in A_{*}(X).$$
\end{proaf}

Using now that the total Milnor class is, up to sign, the difference between the total Schwartz-McPherson and the total virtual class we obtain:

\begin{corollary}\label{11} $${\cal M}(X)=(-1)^{r-1}c\left( \left(
TM|_{X}\right)^{\oplus r-1} \right)^{-1}\cap \displaystyle\sum_{i=1}^{r} a_{1,i}\cdot...\cdot a_{r-1,i} \cdot {\cal M}(X_{i}),$$
where  $a_{j,i}=\left\{\begin{array}{ccl}
c^{Vir}(X_{j+1}) &\;{\rm if}\;& j\leq i\\
c^{SM}(X_{j}) &\;{\rm if}\;& j> i\\
\end{array}\right.$.
\end{corollary}

\begin{corollary} $${\cal M}(X)=(-1)^{n-r} c\left( \left(
TM|_{X}\right)^{\oplus r-1} \right)^{-1}\cap \;\left( c^{Vir}(X_{1})\cdots c^{Vir}(X_{r})- c^{SM}(X_{1})\cdots c^{SM}(X_{r}) \right).$$ \end{corollary}

\section{Applications to line bundles}

In this section we replace the bundles $E_i$ by  line
bundles $L_i$. We assume that each hypersurface
  $X_i$ is equipped with a Whitney stratification  ${\cal S}_{i}$ such that all the intersections amongst strata in the various $X_i$ are transversal.
We set $X := X_1 \cap \cdots \cap X_r$ and we obtain in this section  three applications of the  formulas in the previous section.


\subsection{Parusi\'nski-Pragacz-type formula}
\label{section-P-P-formula}

 Now we  extend to  local complete intersections as above
the following  important characterization of Milnor classes
obtained by A. Parusi\'{n}ski and P. Pragacz in \cite{Par-Pr} for hypersurfaces in compact manifolds:

\begin{equation}\label{P}{\cal M}(X_i):=\sum_{S \in {\cal S}_{i}}\;\gamma_{S}\left(c(L_{i|_{X_i}})^{-1} \cap
c^{SM}(\overline{S})\right)\in A_*(X_i),\end{equation} where
$\gamma_{S}$ is the function defined on each stratum $S$
as follows: For each $x \in S \subset X_i$, let $F_x$ be a
{\it local Milnor fibre} (recall $X_i$ is a hypersurface in the
complex manifold $M$), and  let  $\chi(F_{x})$ be its Euler
characteristic. We set:
$$\mu(x;X_i):= (-1)^{n}\;(\chi(F_{x})-1) \,,$$ and
call it the {\it local Milnor number} of $X_i$ at $x$. This number
is constant on each Whitney stratum, by the topological triviality
of Whitney stratifications, so we denote it $\mu_{S}$. Then
$\gamma_{S}$ is defined inductively by:
$$\gamma_{S}=\mu_{S} - \sum_{S' \neq S,\;\overline{S'} \supset S}
\gamma_{S'}.$$

Following the idea of Parusi\'nski and Pragacz of describing the
Milnor classes using only the Schwartz-MacPherson classes, we have
as a consequence of Corollary \ref{11} and equation (\ref{P}) the
following:

\begin{corollary}\label{ais} The total Milnor class of $X$ is:
$${\cal M}(X)=(-1)^{r-1}c\left( \left(
TM|_{X}\right)^{\oplus r-1} \right)^{-1}\cap \displaystyle\sum_{i=1}^{r}\sum_{S \in {\cal S}_{i}}\;\gamma_{S}\; a_{1,i}\cdot...\cdot a_{r-1,i} \cdot \left(c(L_{i|_{X_i}})^{-1} \cap
c^{SM}(\overline{S})\right),$$
where  $a_{j,i}=\left\{\begin{array}{ccl}
c^{Vir}(X_{j+1}) &\;{\rm if}\;& j\leq i\\
c^{SM}(X_{j}) &\;{\rm if}\;& j> i\\
\end{array}\right.$.\end{corollary}

\vspace{0,2cm} Using \cite[Example 3.2.8 and Proposition 6.3]{Ful}
we have:

\begin{corollary}[Parusi\'{n}ski-Pragacz formula for local
complete intersections]\label{P-P} In the above conditions we have:
 $${\cal M}(X)=(-1)^{nr-n}c\left( \left(
TM|_{X}\right)^{\oplus r-1} \right)^{-1}\cap   \qquad \qquad \qquad
\qquad \qquad \qquad  \qquad \qquad \qquad $$
$$  \qquad \qquad \qquad \qquad \cap \sum{\alpha}_{S_{1},\dots,S_{r}}^{\epsilon_{1}, \dots, \epsilon_{r}}
 \frac{c(L_1)^{\epsilon_{1}}\cdots c(L_r)^{\epsilon_{r}}}{c(L_1\oplus\cdots\oplus L_r)}\cap c^{SM}(\overline{S_1})\cdots c^{SM}(\overline{S_r})\,,$$
 \noindent where the sum runs over all choices  of $S_{i}\in {\cal S}_i,\;i=1,\dots,r$ except $(S_{1},\dots,S_{r})= ((X_1)_{reg},\dots,(X_r)_{reg})$ and
where $(X_{i})_{reg}$ denotes the regular part of $X_{i}$,
 $${\alpha}_{S_{1},\dots,S_{r}}^{\epsilon_{1}, \dots,
\epsilon_{r}}=(-1)^{(n-{1})(\epsilon_{1}+\cdots+\epsilon_{r})}
 \gamma_{S_1}^{1-\epsilon_{1}}\cdots
 \gamma_{S_r}^{1-\epsilon_{r}}\,,$$ $\,$ and

 $$\epsilon_{i}=\left\{\begin{array}{rcl} 1&,& if \;S_{i}=(X_i)_{reg}\\
0&,& if \;\dim(S_i)<n-1\\
\end{array}\right. .$$
\end{corollary}

\subsection{Milnor classes and L\^e cycles}

  Let
${\rm Sing}(X_i)$ be the singular set of $X_i$. This is the set of
points where the section $s_i$ fails to be transversal to the
zero-section of $L_i$. Consider the blow-up $Bl_{\rm Sing(X_i)}M:=B_i$
of $M$ along  ${\rm Sing}(X_i)$, let  $D_i$ be the exceptional divisor
of $B_i$  and ${\cal L}_i$ the associated line bundle on $B_i$, that we
call the {\it tautological line bundle}  of $B_i$. One has a
diagram:
$$
\xymatrix {  &  {\cal L}_i \ar[d]\\
 D_i
\ar[r] \ar[d] & B_i  \ar[d]^{b_i} \\ {\rm Sing}(X_i) \ar[r] & M }
$$

\begin{definition} For $0\leq k \leq n$, we define the $k$-th L\^e cycle of the section
$s_i$ as the following class in the Chow group of $M$:
$$ \Lambda_{k}(X_i)=(b_i)_{*}\left(c_{1}({\cal L}_i)^{n-k} \cap [D_i]\right).$$\end{definition}

\vspace{0.2cm} Notice that $\Lambda_{k}(X_i)$ is supported in
${\rm Sing}(X_i)$, hence we can also think of  $\Lambda_{k}(X_i)$ as being a
class in the Chow group of ${\rm Sing}(X_i)$. Notice also that
$\Lambda_{k}(X_i)$ is zero for $k > {\rm dim}\;{\rm Sing}(X_i)$.

We point out that the cycles obtained in this way are special
cases of {\it Segre cycles} (see for instance \cite{Ful}).
Moreover, because $M$ is compact, these represent classes in the
homology ring $H_{*}(M;\mathbb{Z})$.

In \cite[Theorem 1]{BMS} the authors proved the following
description for the Milnor classes of $X_i$ in terms of the L\^e
cycles of $X_i$.

$${\cal M}_{k}(X_i)=\sum_{l=0}^{d-k}(-1)^{k+l}
\left( \begin{array}{c}l+k\\
k\\
\end{array}\right)
c_{1}(L_i|_{X_i})^{l} \cap \Lambda_{l+k}(X_i) ,$$

Based on this formula and that of Corollary \ref{11}, we get  a
description of Milnor classes of local complete intersections
$X=X_{1} \cap \cdots \cap X_{r}$ via the L\^e cycles of each
hypersurface $ X_{i}$:

\begin{corollary}\label{leformula} For each $0 \le k \le {\rm dim } \,X$, the $k^{th}$ Milnor class of $X$ is:
$${\cal M}_k(X)=(-1)^{r-1}\;c\left( \left(
TM|_{X}\right)^{\oplus r-1} \right)^{-1}\cap \displaystyle\sum_{i=1}^{r} \displaystyle\sum_{l=0}^{d_{i}-k} (-1)^{k+l} \left( \begin{array}{c}l+k\\
k\\
\end{array} \right ) \, \cdot \qquad   \qquad   \qquad   \qquad   \qquad   \qquad  $$
$$  \qquad   \qquad   \qquad   \qquad   \qquad   \qquad   \qquad   \qquad   \qquad  \cdot \, (a_{1,i})\cdot...\cdot (a_{r-1,i}) \cdot\left( c_{1}(L_{i}{|_{X_{i}}})^{l} \cap \Lambda_{l+k}(X_{i})\right) ,$$
where  $(a_{j,i})=\left\{\begin{array}{ccl}
c^{Vir}(X_{j+1}) &\;{\rm if}\;& j\leq i\\
c^{SM}(X_{j}) &\;{\rm if}\;& j> i\\
\end{array}\right.$.\end{corollary}

\subsection{Aluffi type formula}
Let  $X=X_{1} \cap \cdots \cap X_{r}$ as above.
Now we express the total Milnor class of  $X$ in terms of  the Aluffi's
$\mu$-class of each hypersurface $X_i.$ The $\mu$-class was
introduced  in \cite{Aluffi} and it involves the Segre
class of the singular locus of the hypersurface in the total
smooth ambient variety.

For each hypersurface $X_i$ of $M$,  the Aluffi's
 $\mu$-class ${\mu}_{L_i}({\rm Sing}(X_i))$ of the
singular locus of $X_i$  is defined  by the formula
 $$\mu_{L_i}({\rm Sing}(X_i))=c(T^*M\otimes {L_i})\cap s({\rm Sing}(X_i),M),$$ where
$s({\rm Sing}(X_i),M)$ is the Segre class of ${\rm Sing}(X_i)$ in $M$ (see
 \cite[Chapter 4]{Ful}).

We need to introduce some notation.
If $\alpha \in A_*(X_i)$ is a cycle in the Chow group of
$X_i$  and $\alpha=\sum_{j\geq
0}{\alpha}^j,$ where ${\alpha}^j$ is the  codimension $j$
component of $\alpha,$ then Aluffi introduced the following cycles
$$\alpha^{\vee}:=\sum_{j\geq 0}(-1)^j{\alpha}^j \,,$$ and
$$\alpha \otimes L_i:=\sum_{j\geq 0}\frac{{\alpha}^j}{c(L_i)^j}.$$
Then Aluffi proved in \cite{Aluffi} that the total Milnor class
${\cal M}(X_i)$ can be described as follows:

\begin{equation}\label{Aluffi-formula}{\cal M}(X_i)=(-1)^{n-1}c(L_i)^{n-1}\cap (\mu_{L_i}({\rm Sing}(X_i))^{\vee
}\otimes L_i).\end{equation}

\vspace{0,2cm} Again using Corollary \ref{11},  the above
equation yields:
\begin{corollary} \label{Aluffi} The Total Milnor class of  $X := X_1 \cap \cdots \cap X_r$ is:
$${\cal M}(X)=(-1)^{n-r}c\left( \left(
TM|_{X}\right)^{\oplus r-1} \right)^{-1}\cap  \quad \quad\quad\quad  \quad \quad\quad\quad \quad \quad\quad\quad \quad \quad\quad\quad \quad \quad\quad\quad $$  $$  \quad\quad\quad\quad \quad\quad\quad\quad \quad\quad\quad \cap  \left(\displaystyle\sum_{i=1}^{r} a_{1,i}\cdot...\cdot a_{r-1,i}\cdot c(L_{i})^{n-1}\cap (\mu_{L_{i}}({\rm Sing}{(X_{i})\;})^{\vee
}\otimes L_{i})\right),$$
where  $a_{j,i}=\left\{\begin{array}{ccl}
c^{Vir}(X_{j+1}) &\;{\rm if}\;& j\leq i\\
c^{SM}(X_{j}) &\;{\rm if}\;& j> i\\
\end{array}\right.$.
\end{corollary}

\section{General Case}

\hspace{0.5cm} Let $M$ be an $m$-dimensional compact complex
analytic manifold and let $E$ be a holomorphic vector bundle over
$M$ of rank $r$. Let $Z(s)$ be the zero set of a regular holomorphic
section $s$ of $E$, so this is an $(m-r)$-dimensional local complete
intersection.

Let $p:N\rightarrow M$ be a proper morphism, where $N$ is an $n$-dimensional compact complex
analytic manifold. Consider the diagram
$$
\xymatrix {  \mathbb{P}(T^{*}N)  \ar[dr]^{\pi_{N}} & \mathbb{P}(p^{*}(T^{*}M)) \ar[l]_{i}
\ar[r]^{g} \ar[d]^{p^{*}(\pi_{M})} & \mathbb{P}(T^{*}M) \ar[d]^{\pi_{M}}  \\ & N \ar[r]^{p}  & M  }
$$
where $\pi_{M}:{\mathbb P}(T^{*}M)\rightarrow M$  and $\pi_{N}:{\mathbb P}(T^{*}N)\rightarrow N$ are the corresponding projectivized cotangent bundles.

\begin{lemma}\label{M4}  Let $Ch({\mathbbm 1}_{Z(s)})$ be the characteristic cycle of the characteristic function ${\mathbbm 1}_{Z(s)}$. Then $$i_{*}g^{*}\;[\mathbb{P}(Ch({\mathbbm 1}_{Z(s)}))]=(-1)^{n-m}\;
[\mathbb{P}(Ch({\mathbbm 1}_{Z(p^{*}s)}))].$$ \end{lemma}

\begin{proof} Let $\{S_{\alpha}\}$ be a Whitney stratification of $Z(s)$. Then
$\{p^{-1}(S_{\alpha})\}$ is a Whitney stratification of
$Z(p^{*}s)$. By definition, $$\mathbb{P}(Ch({\mathbbm 1}_{Z(s)}))=\sum
m_{\alpha}\mathbb{P}\left(\overline{T_{S_{\alpha}}^{*}M}\right),$$ where
$m_{\alpha}:= (-1)^{m-r-1}\chi\left(\phi_{h_{1}|Z(s)}H_{1}^{\bullet}
\right)_{z_{1}}$ with $z_{1} \in S_{\alpha}$, $H_{1}^{\bullet}$ is the complex of sheaves
such that  $\chi( H_{1}^{\bullet})_{z}={\mathbbm 1}_{Z(s)}(z)$,  $h_{1}:
(M,z_{1})\rightarrow (\C,0)$ is a germ  that  satisfies $(z,d_{z}h_{1}) \in
\overline{T_{S_{\alpha}}^{*}M}$ and  $(z,d_{z}h_{1}) \not\in
\overline{T_{S_{\beta}}^{*}M}$ for all strata $S_{\beta} \neq S_{\alpha}$ and $\phi_{h_{1}|Z(s)}H_{1}^{\bullet}$ is the sheaf of vanishing cycles of $h_{1}$ restricted to $Z(s)$.
Analogously,
$$\mathbb{P}(Ch({\mathbbm 1}_{Z(p^{*}s)}))=\sum
n_{\alpha}\mathbb{P}\left(\overline{T_{p^{-1}(S_{\alpha})}^{*}N}\right),$$
where $n_{\alpha}:=
(-1)^{n-r-1}\chi\left(\phi_{h_{2}|Z(p^{*}s)}H_{2}^{\bullet}
\right)_{z_{2}}$, with $z_{2} \in p^{-1}(S_{\alpha})$, $H_{2}^{\bullet}$ is the
complex of sheaves such that   $\chi(
H_{2}^{\bullet})_{q}={\mathbbm 1}_{Z(p^{*}s)}(q)$, $h_{2}:
(N,z_{2})\rightarrow (\C,0)$ is a germ  that satisfies $(z_{2},d_{z_{2}}h_{2}) \in
\overline{T_{p^{-1}(S_{\alpha})}^{*}N}$ and  $(z_{2},d_{z_{2}}h_{2}) \not\in
\overline{T_{p^{-1}(S_{\beta})}^{*}N}$ for all strata $S_{\beta} \neq
S_{\alpha}$.

\vspace{0.2cm}These definitions do not depend on the choices of $
(z_{1},h_{1}) $ and $(z_{2},h_{2}) $ respectively. So we first  choose $ z_{1} $ and $ h_{1} $ in the
first definition,   then fix a $z_2 \in p^{-1}(z_1)$ and set   $ h_{2}=h_{1} \circ p $.

Note that $$p^{*}H_{1}^{\bullet}=H_{2}^{\bullet} \quad \hbox { and } \quad
\phi_{p^{*}(h_{1}|Z(s))}p^{*}H_{1}^{\bullet}=p^{*}\left(\phi_{h_{1}|Z(s)}H_{1}^{\bullet}\right)\,.$$
Thus we have that $\chi\left(\phi_{h_{2}|Z(p^{*}s)}H_{2}^{\bullet}
\right)_{z_{2}}=\chi\left(\phi_{h_{1}|Z(s)}H_{1}^{\bullet}
\right)_{z_{1}}$. Hence
\begin{equation}\label{M6}m_{\alpha}=(-1)^{n-m} n_{\alpha}. \end{equation}

Note that
\begin{equation}\label{M7}i_{*}g^{*}\;[\mathbb{P}\left(\overline{T_{S_{\alpha}}^{*}M}\right)]=
[\mathbb{P}\left(\overline{T_{p^{-1}(S_{\alpha})}^{*}N}\right)].\end{equation}
Therefore, by equations (\ref{M6}) and (\ref{M7}) we get,
$$i_{*}g^{*}\;[\mathbb{P}(Ch({\mathbbm 1}_{Z(s)}))]=(-1)^{n-m}\;
[\mathbb{P}(Ch({\mathbbm 1}_{Z(p^{*}s)}))],$$
as stated.
\end{proof}

\begin{lemma}\label{1} In the conditions above, if  the morphism $p:N\rightarrow M$ is also flat, then  $$p^{*}\left({\cal M}(Z(s))\right)=\displaystyle\frac{c(p^{*}(TM))}{c(TN)}\cap {\cal M}(Z(p^{*}s)).$$
\end{lemma}

\begin{proof} Since $p$ is flat, $p^{*}[Z(s)]=[Z(p^{*}s)]$ (by \cite[Proposition 14.1]{Ful}).
Hence,
$$p^{*} c^{Vir}(Z(s))=p^{*}(c(TM).c(E)^{-1} \cap [Z(s)])=c(p^{*}TM) c(p^{*}E)^{-1} \cap p^{*}[Z(s)] \quad$$
$$\qquad \quad \qquad  \qquad =\frac{c(p^{*}TM)}{c(TN)}
c(TN) c(p^{*}E)^{-1} \cap [Z(p^{*}s)]=\frac{c(p^{*}TM)}{c(TN)}\cap c^{Vir}(Z(p^{*}s)).$$
On the other hand, using the description of Schwartz-MacPherson classes due to C.
Sabbah in \cite{Sabbah}, we have:
$$p^{*}c^{SM}(Z(s))=(-1)^{m-1}\frac{c(p^{*}TM)}{c(TN)}c(TN)\cap
p^{*}\left({\pi_{M}}_* \left( c({\cal O}_{M}(1))^{-1}\cap
[\mathbb{P}(Ch({\mathbbm 1}_{Z(s)}))]\right)\right),$$ where
${\cal O}_{M}(1)$ denotes the tautological line bundle on the
projectivized cotangent bundle $\pi_{M}:{\mathbb
P}(T^{*}M)\rightarrow M$. By Lemma \ref{M4}, we have: {\small
$$p^{*}\left({\pi_{M}}_* \left( c({\cal O}_{M}(1))^{-1}\cap
[\mathbb{P}(Ch({\mathbbm 1}_{Z(s)}))]\right)\right)=
(-1)^{n+m}{\pi_{N}}_* \left( c({\cal O}_{N}(1))^{-1}\cap
[\mathbb{P}(Ch({\mathbbm 1}_{Z(p^{*}s)}))]\right),$$} where ${\cal
O}_{N}(1)$ is the tautological line bundle on the projectivisation
$\pi_{N}:P(T^{*}N)\rightarrow N$. Hence,
$$p^{*}c^{SM}(Z(s))=\frac{c(p^{*}TM)}{c(TN)}\cap \left((-1)^{n-1} c(TN)\cap
{\pi_{N}}_* \left( c({\cal O}_{N}(1))^{-1}\cap
[\mathbb{P}(Ch({\mathbbm 1}_{Z(p^{*}s)}))]\right)\right)$$
$$=\frac{c(p^{*}TM)}{c(TN)}\cap c^{SM}(Z(p^{*}s)). \qquad  \quad  \qquad  \qquad   \qquad  \qquad  \qquad $$
Therefore: $p^{*}\left({\cal M}(Z(s))\right)=\displaystyle\frac{c(p^{*}(TM))}{c(TN)}\cap {\cal M}(Z(p^{*}s)).$
\end{proof}

Now we need:.

\begin{lemma} \label{2} Consider  an exact sequence of vector bundles on $M$:
$$0\longrightarrow F \longrightarrow G\longrightarrow H\longrightarrow 0 \,.$$
Given a section $t$ of $G$,  let $t_H$ be the induced section of $H$. Set $Z_{1}=Z(t_{H})$ and let $\tilde t$ be the  section of  $F|_{Z_1}$
induced by $t|_{Z_1}$. Then, $${\cal M}(Z(t))=c(F)^{-1} c_{top}(F) \cap {\cal M}(Z_{1})\,,$$ where $c_{top}(F)$ is the top Chern class of $F$.
\end{lemma}

\begin{proof} Using \cite[Proposition 1.3]{PP}  we have:
$$c^{SM}(Z(t))=c(F)^{-1}c_{top}(F)\cap c^{SM}(Z_{1}).$$
On the other hand, let us consider the following virtual bundles:

$$\tau (Z_{1},M)=TM|_{Z_{1}}-H|_{Z_{1}} \,,$$
$$\tau (Z(t),M)=TM|_{Z(t)}-G|_{Z(t)} \,,$$
$$\quad \tau (Z(t),Z_{1}):=\tau (Z_{1},M)-i^{*}F|_{Z(t)} \,.$$
Notice that $c_{top}(i^{*}F) \cap [Z_{1}]=j_{*}[Z(\tilde t)]$ where $j: Z(\tilde t) \to Z_1$ is the inclusion. On the other hand $Z(\tilde t) = Z(t)$.
Hence we have:
$$c_{top}(i^{*}F) \cap c^{Vir}(Z_{1},M)=c_{top}(i^{*}F) \cap c(TM).c(H)^{-1}\cap [Z_{1}] \,$$  $$\qquad \quad \qquad \qquad =c(TM).c(H)^{-1}\cap j_{*}[Z(t)].$$
Then, setting  $c^{Vir}(Z(t),Z_{1}) := c({\tau}(Z(t),Z_{1}))$, we get:
$$c(F)^{-1}.c_{top}(F) \cap c^{Vir}(Z_{1},M)=c^{Vir}(Z(t),Z_{1})=c^{Vir}(Z(t),M),$$
{\it i.e.}, $c^{Vir}(Z(t))=c(F)^{-1}.c_{top}(F) \cap c^{Vir}(Z_{1})$. Thence: $${\cal M}(Z(t))=c(F)^{-1} c_{top}(F) \cap {\cal M}(Z_{1})\,,$$ as stated.
\end{proof}

\begin{theorem}\label{t:final} Let $s: M \to E$ be a regular section of the holomorphic bundle $E$. Let
$p:\mathbb{P}(E^{\vee})\rightarrow M$ be the projectivization of the dual bundle $E^{\vee}$, so we have the tautological exact sequence   $0\rightarrow F\rightarrow p^{*}E\rightarrow {\cal O}_{\mathbb{P}(E^{\vee})}(1)\rightarrow 0$. Then,
$${\cal M}(Z(s)) = p_{*} \Big(\big[c\big(p^{*}(E^{\vee}) \otimes {\cal O}_{{\mathbb P}(E^{\vee})}
(1)\big)^{-1}\cdot\,
 c_{1}\big({\cal O}_{\mathbb{P}(E^{\vee})}(1)\big)^{r-1}\cdot\, c(F)^{-1}\cdot\, c_{top}(F)\big] \cap {\cal M}(Z(\tilde{s}))\Big),$$
where $\tilde{s}$ is the section induced by $p^{*}s$ in ${\cal
O}_{\mathbb{P}(E^{\vee})}(1)$.

 \end{theorem}

 \begin{proof} By the Lemma \ref{1}, $$p^{*}\left({\cal M}(Z(s))\right)=\displaystyle\frac{c(p^{*}TM)}{c(T\mathbb{P}(E^{\vee}))}\cap {\cal M}(Z(p^{*}s)).$$
It follows from \cite[Proposition 3.1]{Ful} that
$${\cal M}(Z(s))=p_{*}\left( c_{1}({\cal O}_{\mathbb{P}(E^{\vee})}(1))^{r-1}\frac{c(p^{*}TM)}{c(T\mathbb{P}(E^{\vee}))} \cap{\cal M}(Z(p^{*}s)) \right).$$
Using  Lemma \ref{2}
we have: $${\cal M}(Z(p^{*}s))=c(F)^{-1}c_{top}(F)\cap {\cal M}(Z(\tilde{s}))\,,$$ and we arrive to
\begin{equation}\label{tangsequence0} {\cal M}(Z(s)) = p_{*} \left( c_{1}({\cal O}_{\mathbb{P}(E^{\vee})}(1))^{r-1} \frac{c(p^{*}TM)}{c(T\mathbb{P}(E^{\vee}))}\;c(F)^{-1}c_{top}(F) \cap {\cal M}(Z(\tilde{s}))\right).\end{equation}
Now using the exact sequence  $$0\rightarrow T_{{\mathbb P}(E^{\vee})/{M}}\rightarrow T{{\mathbb P}(E^{\vee})}\rightarrow p^{*}(TM)\rightarrow 0 \,,$$
we get
\begin{equation}\label{tangsequence1} \frac{c(p^{*}(TM))}{c(T{{\mathbb P}(E^{\vee})})} = c(T_{{\mathbb P}(E^{\vee})/{M}})^{-1} \,.\end{equation}
On the other hand, using the exact sequence  $$0\rightarrow {\cal O}_{{\mathbb P}(E^{\vee})}\rightarrow p^{*}(E^{\vee}) \otimes {\cal O}_{{\mathbb P}(E^{\vee})} (1)\rightarrow T_{{\mathbb P}(E^{\vee})/{M}}\rightarrow 0 \,,$$
we get
\begin{equation}\label{tangsequence2} c(p^{*}(E^{\vee}) \otimes {\cal O}_{{\mathbb P}(E^{\vee})} (1))=c(T_{{\mathbb P}(E^{\vee})/{M}})\end{equation}
Therefore by equations (\ref{tangsequence0}), (\ref{tangsequence1}) and (\ref{tangsequence2}) we obtain:
$${\cal M}(Z(s)) = p_{*} \Big(\big[c\big(p^{*}(E^{\vee}) \otimes {\cal O}_{{\mathbb P}(E^{\vee})} (1)\big)^{-1} c_{1}\big({\cal O}_{\mathbb{P}(E^{\vee})}(1)\big)^{r-1} c(F)^{-1}c_{top}(F)\big] \cap {\cal
M}(Z(\tilde{s}))\Big).$$
 \end{proof}

 \end{document}